### v2

# Cascade-driven opinion dynamics on social networks

Abstract
Online social networks (OSNs) have transformed the way individuals fulfill their social needs and consume information. As OSNs become increasingly prominent sources for news dissemination, individuals often encounter content that influences their opinions through both direct interactions and broader network dynamics. In this paper, we propose the Friedkin-Johnsen on Cascade (FJC) model, which, to the best of our knowledge, is the first attempt to integrate information cascades and opinion dynamics, specifically using the very popular Friedkin-Johnsen model. Our model, validated over real social cascades, highlights how the convergence of socialization and sharing news on these platforms can disrupt opinion evolution dynamics typically observed in offline settings. Our findings demonstrate that these cascades can amplify the influence of central opinion leaders, making them more resistant to divergent viewpoints, even when challenged by a critical mass of dissenting opinions. This research underscores the importance of understanding the interplay between social dynamics and information flow in shaping public discourse in the digital age.
Index Terms:
Information cascades, opinion dynamics, polarization, Friedkin-Johnsen model, viral content
I Introduction
Online social networks (OSNs) have fundamentally transformed how individuals communicate, share information, and form opinions in the digital age. Serving both as platforms for social interaction and as primary sources of real-time news, OSNs facilitate unprecedented levels of connectivity and information dissemination [1]. Additionally, OSNs have emerged as prominent sources of real-time news dissemination, with a substantial proportion of the population relying on these platforms for information [1]. Political figures, journalists, and ordinary users frequently utilize OSNs to disseminate announcements and opinions, rendering them integral to contemporary information consumption patterns. Moreover, media organizations have implemented strategies to encourage the sharing of their content on OSNs, facilitating the seamless integration of news into users’ social interactions. Therefore, during our social activities online, we can also be exposed to the news even incidentally. However, this convergence of socialization and news consumption on OSNs has raised concerns regarding the formation of filter bubbles and echo chambers, wherein individuals are exposed primarily to content that reinforces their existing beliefs. Such environments pose risks of fostering biased perspectives and perpetuating misinformation, particularly with the proliferation of fake news campaigns designed to manipulate public opinion [2]. These dynamics underscore the complex interplay between socialization, news consumption, and information dissemination in the digital era, warranting further empirical investigation and theoretical exploration within interdisciplinary frameworks. Understanding the mechanisms behind these phenomena requires an integrated study of how opinions evolve through social influence and how information cascades propagate within these networks.
The dynamics of opinion formation represent a multifaceted research area situated at the intersection of various disciplines, including sociology, economics, psychology, control theory, computer science, computational social science, and physics. Numerous models have been proposed to describe the mechanisms underlying opinion dynamics, aiming to capture observed theoretical and empirical phenomena. These models often account for factors such as individuals’ inclination to reduce cognitive dissonance with close acquaintances or their resistance to relinquishing deeply held biases. Among the prominent models in this domain is the Friedkin-Johnsen (FJ) model [3] belonging to the category of averaging models [4, 5, 3], according to which people iteratively adjust their opinions towards the average opinion of their neighbours, with the strength of social ties weighting the influence of each neighbor’s opinion. Additionally, the FJ model incorporates a parameter representing individuals’ open-mindedness, referred to as susceptibility, which modulates the extent to which individuals revise their opinions. The FJ model has gained popularity due to its mathematical tractability and its empirical validation on real social networks. It has demonstrated the ability to replicate various observed mechanisms in social networks, including the phenomenon of wisdom of the crowd [6] and the emergence of polarization [7] and recently, it has also been proposed as the building block to model multilateral international negotiation processes [8].
The Friedkin-Johnsen (FJ) model has been validated in offline settings, where opinion evolution reflects direct interpersonal interactions and deliberations across various topics [9, 10]. However, online social environments introduce distinct mechanisms of opinion exchange: individuals are exposed to their contacts’ expressed opinions indirectly, primarily through public social media activities such as posts, comments, and shares [11, 12]. This mode of exposure frequently involves repeated encounters with viewpoints from a broader and more heterogeneous audience, which can heighten cognitive dissonance by presenting conflicting information [13, 14]. In online environments, opinion exposure is both asynchronous and large-scale, allowing conflicting signals to endure over time and reach diverse audiences. At the same time, empirical studies show that controversial or polarizing opinions propagate with exceptional speed and breadth compared to offline interactions, amplifying their influence on collective attitudes [15, 16]. These opinion expressions propagate through various content formats, including textual posts, news resharing, and multimedia, whose spread dynamics are effectively modeled by the theory of information cascades [17].
Opinion dynamics and information cascades operate at distinct granularity and temporal scales. Information cascades focus on how individual pieces of information (e.g., a post, an image) are shared and propagate within the social network, while opinion dynamics concern the evolution of opinions through interactions among individuals. Both processes exhibit emergent behaviors, transitioning from local interactions to global effects such as content virality and polarization. In opinion dynamics, the specific social interactions leading to opinion change remain abstract and unspecified. Conversely, information cascades emphasize the sharing mechanism of content without addressing the effects of such sharing. Importantly, reposting online content received from friends (e.g., retweets) is one of the most common forms of interaction in digital platforms. Empirical evidence shows that in a complete 24-hour snapshot of Twitter activity, only 20.8% of tweets were original, while 50.7% were retweets, underscoring that resharing dominates user behavior [18]. This widespread resharing suggests a natural intertwining between opinion dynamics and information cascades: as users repeatedly encounter and propagate content, microscopic acts of resharing can influence macroscopic opinion evolution. Despite this evident overlap, the two processes have thus far been treated separately in the literature. In this work, we set out to fill this gap and establish an integrated model whereby the microscopic act of content sharing bears an effect in the macroscopic phenomenon of opinion evolution in online social networks.
To address this issue, we propose the Friedkin-Johnsen on Cascade (FJC), a novel model that integrates the FJ opinion dynamics framework with information cascades. Specifically, we hypothesize that opinions on a selected topic are expressed and disseminated through the same flow that governs the resharing of online posts. A post is initially shared by certain sources and then propagates as a cascade through successive neighborhoods until it loses relevance or reaches saturation. When a node reshapes a post (e.g., via a retweet with a comment or framing), it not only amplifies the content but also attaches its own expressed opinion on the topic—such as a textual remark, endorsement, or critique—thereby exposing neighbors to both the content and the sharer’s stance. Receiving such a reposted message from a neighbor can trigger an opinion update in the recipient, linking the micro-level cascade of reposting to the macro-level evolution of opinions.
First, we validated our FJC model on a real dataset of information cascades from Twitter, demonstrating its strong predictive power for online opinion dynamics. Building upon our previous work [7], we then analyzed the impact of cascade mechanisms on polarization, evaluating polarization levels using opinion vectors that maximize polarization in the classical FJ model. To this end, we tested the model on both synthetic and real social networks. Our main findings can be summarized as follows:
  • Highly central nodes emerge as opinion leaders with disproportionate influence, capable of either intensifying or mitigating polarization. Notably, regardless of the extent of information spread, the sequential nature of cascades causes central nodes to encounter content early, limiting their exposure to diverse opinions and making them less susceptible to external influence.
  • The FJC dynamics weaken the influence of peripheral nodes—even when they are numerous, stubborn, and aligned—since they often fail to reach and sway the central nodes.
  • Information cascades can substantially shape polarization outcomes, either amplifying or reducing them, depending on how opinions spread through the network.
These findings underscore how sharing mechanisms on online platforms introduce structural and temporal asymmetries that reshape opinion evolution and amplify the role of network topology compared to the classical FJ model. Furthermore, they highlight how online dynamics can exacerbate polarization beyond what is typically observed in offline social contexts. Neglecting the role of diffusion dynamics can therefore lead to serious misjudgments in assessing both opinion formation and polarization. These insights highlight the critical role of online information diffusion in shaping collective attitudes and social interactions and offer guidance for communication strategies in business, public discourse, and policy-making.
II Related Works
The modeling of opinion dynamics has garnered the interest of researchers for decades, with this interest further amplified with the emergence of online social networks. There are several classes of opinion dynamics models, and the literature on the topic is extensive [19]. Here, we concentrate on the class of averaging models, which posit that users adjust their opinions based on those of the people around them to mitigate social discomfort. Among these models, the FJ model [3] is one of the most studied [20, 21, 22, 23, 24, 25, 26, 10, 27, 7]. Its inner workings are explained in Section III-C. Recently, research on the FJ model has taken two main directions. The first involves optimization problems based on FJ [28, 29, 30, 23, 20, 21], whose main goal is to make interventions into the social network to achieve specific desired outcomes in the final opinions. For example, some investigate how to modify the initial opinion of a set of nodes to maximise the opinion discord [28] or minimize the overall opinion [29]. The second direction explores the possible extensions of the FJ model to capture diverse observed features in the online social networks [31, 32, 33, 34, 35]. In particular, these papers incorporate both attractive and repulsive relationships between users [32, 34], even assuming that the influence of a node on another can vary according to the homophily principle [35]. Furthermore, Bhalla et al. in [31] analyzes the polarization effect of FJ on time-varying networks and Chitra et al. in [33] extended the FJ to reproduce filter bubbles. Differently from these works, we investigate the effect of the classical FJ model on a static network when the social debate is shaped by the temporal flow of news posts on online social platforms.
To model the posting flow, we mainly take hints from the very popular independent-cascade model in [36] (see Section III-B for a description of how it works), which is able to reproduce realistic diffusion dynamics [37]. This seminal work has inspired many following studies and extensions [38, 39, 40, 41], which extend the classical model to include collaborative and/or antagonistic relationships among users, as well as dynamic networks. The main goal of information diffusion is to study the rate of diffusion of viral content depending on the parameters. Instead, we are not interested in the popularity of content, but rather in its diffusion process. Hence, we used the original independent-cascade model and analyzed its impact on the opinion formation process. Similarly to opinion dynamics, the literature on information diffusion is extensive, so we refer the interested reader to dedicated surveys for more details [42].
The works most closely related to ours are [43, 44], which also investigate the interplay between information diffusion and opinion formation. While [43] adopts the Friedkin–Johnsen (FJ) model and [44] employs a bounded confidence framework, both approaches differ fundamentally from ours. In [43], content influences opinions only once, by shifting initial biases based on content popularity; subsequent opinion dynamics unfold independently of further information diffusion. In [44], opinion adoption is driven solely by similarity to the root node’s opinion and is primarily used to model content virality, not the cognitive process of opinion change. In contrast, our FJC model assumes that every act of content sharing conveys the sharer’s current opinion, which in turn directly influences the recipient’s belief. This enables us to explicitly capture how the temporal structure of cascades shapes long-term opinion dynamics and polarization—an aspect not addressed in prior work.
III Preliminaries
In this section, we will summarize the theoretical notions and notation we will use throughout the rest of the paper.
III-A The social network
We can describe an online social network platform as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where the $v$ nodes $\mathcal{V}$ represent users and the edges $\mathcal{E}$ represent social connections between them. In the most general form, the graph $G$ is directed and the edges $(i,j)$ represent a follow relationship. Specifically, $(i,j)\in\mathcal{E}$ means that $i$ follows $j$, thus $j$ influences $i$ as its activity appears in $i$’s news feed. The set of nodes that has an influence on $i$ is thus composed of $i$’s out-neighbours, which we will denote with $\mathcal{N}^{+}(i)$. Instead, the set of nodes that is influenced by node $j$ is the in-neighbourhood of $j$, which we indicate with $\mathcal{N}^{-}(j)$ [^1] Please note that, generally, the neighbourhood of a node corresponds to our out-neighbourhood.. It is worth noting that in the special case of an undirected node, $\mathcal{N}^{+}(i)$ and $\mathcal{N}^{-}(i)$ are the same. This social connection can arise, for example, when $i$ follows $j$ on platforms like X or Facebook or when they are friends on Facebook. The strength of this relationship is mathematically represented by the quantity $\hat{w}_{ij}$, which is the $(i,j)$-th entry of the $n\times n$ matrix $\hat{W}$ that we call social matrix. The weight that node $i$ gives to its out-neighbours represents the degree of influence of their opinions. It is provided by the $(i,j)$-th entry of the $n\times n$ matrix $W$, named the influence matrix $W$, which is row-stochastic. On the $i$-th row, for example, we can see how the influence on $i$’s opinion is distributed among all nodes in $i$’s network. In situations where additional data is lacking, the matrix $W$ can be derived from $\hat{W}$ through row-normalization, expressed as $w_{ij}=\frac{\hat{w}{ij}}{\sum_{j}\hat{w}{ij}}$. This process capitalizes on the insight that stronger relationships should carry greater weight in opinion formation. It’s worth noting that, with this mechanism, even in cases where the social graph exhibits symmetry, the influence matrix results asymmetric.
III-B Information cascade
Our model conceptualizes the act of posting social content following the independent cascade model [36]. In the independent cascade model, each node can be active or inactive. An initial set of active nodes $v$ has a single chance to activate their neighbours $w$, and each activation is successful with probability $p_{vw}$. The nodes $w$ that are successfully activated become active at the next step and try to activate their neighbours as well. The process continues until no other nodes can be activated. This framework is well-suited to modeling the propagation of shared elements in online social networks, such as tweets or posts, where each user typically has one opportunity to reshare content upon seeing it, and further reshares depend on their direct followers.
III-C Friedkin-Johnsen opinion dynamic model
The FJ opinion dynamic model is a mathematical model that describes how people’s opinions change over time due to their mutual influences. It does this by representing people’s opinions as real numbers and then modelling two opposing forces that affect how those opinions evolve. The first force is an initial bias that each person $i$ has, which is represented by $u_{i}$. The second force is a desire to reduce the discomfort people feel when their opinions differ from those around them, as described by the cognitive dissonance theory. Each person $i$ has a parameter $\lambda_{i}$, called susceptibility, that determines how much weight they give to each of these forces. If $\lambda_{i}$ is 0, it means the person is completely stubborn and will never change their mind. If $\lambda_{i}$ is 1, it means the person is completely naive and will totally get over their initial bias. Node susceptibilities are generally represented by the diagonal matrix $\Lambda$, whose $i$-th diagonal element is equal to the susceptibility $\lambda_{i}$.
The FJ model is mathematically expressed as a simple update equation that is applied to each person’s opinion at discrete times. As discussed in our previous work [7], in the literature, the term “FJ model” refers to many variants of the original model. The update expression for the most general variant of the FJ model is given by the following equation:
$x_{i}(t+1)=\lambda_{i}\sum_{j=0}^{v}w_{ij}x_{j}(t)+(1-\lambda_{i})u_{i}\qquad\forall i,$ (1)
where $w_{ij}$ is the $(i,j)$-th entry of matrix $W$. The first term of the sum represents the movement of $i$’s opinion towards the opinion of its neighbours, weighted by their influence degree. The second term represents instead the anchoring to the initial prejudice. Under general assumptions, the model converges, as summarized in the following theorem (proved in [25]).
Theorem 1.
[FJ convergence] If the graph is strongly connected and $\Lambda\neq I$ (i.e., at least one agent partially anchored to its initial opinion exists) [^2] Please note that when $\Lambda=I$, the FJ model reduces to the French-DeGroot model [4], which is a simpler model that has been separately studied in the literature. then the FJ model is convergent, and it converges to the stable opinion vector given by:
$\bm{z}=(I-\Lambda W)^{-1}(I-\Lambda)\bm{u},$ (2)
where $I$ is the identity matrix.
This version of the FJ model is synchronous, meaning that all nodes update their opinion simultaneously. However, in reality, people engage in asynchronous interactions, so opinions do not evolve simultaneously. In other words, we do not engage with all our social contacts all at once. To capture this, the FJ model also has an asynchronous version [45] that, as we will discuss later, has been proven to converge to the synchronous FJ model of Equation (1). In the asynchronous model, an edge $(i,j)$ is selected uniformly at random at each time step and the opinion of only one of the two nodes (say $i$, selected at random) is updated, according to the following equation
$x_{i}(t+1)=h_{i}\left[(1-\gamma_{ij})x_{i}(t)+\gamma_{ij}x_{j}(t)\right]+(1-h_{i})u_{i},$ (3)
while the rest of the nodes maintain their opinion. Note that the term $(1-\gamma_{ij})x_{i}(t)$ balances the current expressed opinion of node $i$ against that of its neighbor $j$, with $\gamma_{ij}$ controlling the strength of influence from $j$. Here matrices $H=\{h_{ij}\}$ and $\Gamma=\{\gamma_{ij}\}$ replace $\Lambda$ and $W$ of the synchronous model, and are different from them. In fact, to guarantee the convergence of the model in Equation (3) to that in Equation (1), the formulation has to take into account the probability of node $i$ being randomly selected, which, in turn, depends on the degree of nodes. Since high-degree nodes are involved in more edges, when selecting edges at random high-degree nodes have a higher probability of being selected. For this reason, $H$ and $\Gamma$ should be defined from $\Lambda$ and $W$ as follows:
$h_{i}=1-\frac{1-\lambda_{i}}{d_{i}}$ (4)
$\gamma_{ij}=\left\{\begin{array}[]{ll}\frac{d_{i}(1-h_{i})+h_{i}-(1-\lambda_{i}w_{ii})}{h_{i}}&\textrm{if }i=j\\ \frac{\lambda_{i}w_{i}j}{h_{i}}&\textrm{if }i\neq j\end{array}\right$. (7)
where $d_{i}$ is the degree of node $i$. In [45], Frasca et al. establishes the ergodic convergence of the asynchronous FJ algorithm defined by equations (3) to the classical FJ. Notably, while the individual state variables generally do not converge in the pointwise sense and often exhibit persistent oscillation, their time-averaged values converge asymptotically to a stable limit, which, in our case, coincides with the stable point of the FJ.
III-D Polarization
In our previous work [7], we reviewed and classified the most popular polarization metrics used in the related literature. These metrics, given an opinion vector, measure polarization according to different semantics. They can be summarized in the following definition.
Definition 1.
For an opinion $\bm{x}=(x_{i})\in[-1,1]^{v}$ the following polarization indices are defined:
$NDI(\bm{x})$ $=\sum_{(i,j)\in\mathcal{E}}w_{ij}(x_{i}-x_{j})^{2}$ (8)
$GDI(\bm{x})$ $=\sum_{i,j\in\mathcal{V}:i<j}(x_{i}-x_{j})^{2}$ (9)
$P_{1}(\bm{x})$ $=\sum_{i\in\mathcal{V}}(x_{i}-\bar{\bm{x}})^{2}=\|\bm{x}-\bar{\bm{x}}\|_{2}^{2}$ (10)
$P_{2}(\bm{x})$ $=\frac{1}{|\mathcal{V}|}\sum_{i\in\mathcal{V}}x_{i}^{2}=\frac{1}{|\mathcal{V}|}\|\bm{x}\|_{2}^{2}$ (11)
$P_{3}(\bm{x})$ $=\sum_{i\in\mathcal{V}}x_{i}^{2}=\|\bm{x}\|_{2}^{2}$ (12)
$P_{4}(\bm{x})$ $=\sum_{i\in\mathcal{V}}|x_{i}|=\|\bm{x}\|_{1}$ (13)
In the context of measuring polarization in social networks, the metrics $NDI$, GDIGDI, and $P_{1}$ quantify polarization based on the degree of disagreement among individuals’ opinions. On the other hand, metrics $P_{2}$, $P_{3}$, and $P_{4}$ consider polarization to occur when opinions are significantly distant from a neutral opinion, which is assigned a value of 0.
A dynamic opinion model is considered polarizing with respect to a given metric $\Phi$ if there exists at least one initial opinion vector $\bm{u}$ such that the corresponding final opinion $\bm{z}$ exhibits greater polarization, i.e., $\Phi(\bm{z})>\Phi(\bm{u})$. The quantity
$\Delta_{\Phi}(\bm{u})=\Phi(\bm{z})-\Phi(\bm{u})$ (14)
is referred to as the polarization value and, by definition, is positive if and only if the model is polarizing with respect to the metric $\Phi$. Our previous research [7] showed that the FJ model is inherently polarizing and provided formulas to derive polarizing vectors for the polarization metrics $P_{2}$, $P_{3}$, and $P_{4}$. These vectors are closely linked to the spectral properties of the matrix $H=(I-\Lambda W)^{-1}(I-\Lambda)$, which maps initial opinions $\bm{u}$ to final opinions $\bm{z}$ as in Equation (2). Depending on the considered sub-domain of the opinion space, different polarizing vectors can be identified. In this paper, for $P_{4}$ polarization we use $\bm{u}_{B_{1}(1)}$, while for $P_{2}$ and $P_{3}$ polarization we rely on $\bm{u}_{B_{2}(1)}$, $\bm{u}_{B_{2}(t)}$, and the heuristic $\bm{u}_{V_{>1}}^{heu}$, listed in order of increasing polarization strength. For a detailed mathematical treatment and derivations, we refer the reader to [7]. In Section V we will examine the impact of information cascades when the opinion dynamic process starts with users featuring these polarizing opinions. Notably, we will observe that FJC model induces significantly greater polarization than the FJ for most of the scenarios studied.
IV The Friedkin-Johnsen on Cascades model
In this section, we will discuss the FJC (Friedkin-Johnsen on Cascades) model. The FJ model is an opinion dynamic model that explains how people form their opinions in the offline world, where face-to-face interactions significantly influence the depth of relationships and consequently the shared understanding of opinions. This is the unique mechanism that triggers the tension resolving process, which is the brain’s reaction to cognitive dissonance. Information cascades, on the other hand, describe how information is spread through online social networks, where interactions are encoded as engagement with each other’s posts. They characterise how a single post by a source can go viral and inundate the entire network. Despite belonging to two different and separate research areas, opinion dynamics and information cascades are intricately linked. In online social networks, opinions are expressed in a cascade manner, by subsequent exposure to other contents/opinions. It is reasonable to expect that this unique exposure to others’ posts influences how we update our opinions. The amplification of virality in social networks can disrupt the balance of real-life discussions in offline networks, thus strengthening the role of central nodes. This is a crucial insight that has far-reaching implications for understanding opinion formation and information dissemination in online social networks.
IV-A Revisiting the independent cascade model
The standard independent cascade model needs to be slightly adjusted to integrate it into the FJC framework. To this aim, we identify three new node states, which replace the previous ones: unaware, aware and spreader. A node is unaware when it has not been exposed to the post yet. As soon as it “sees” the post from a neighbour, the node becomes aware and then it updates its opinion as described in the following sections. With a probability $\theta_{s,i}$ it also reshares the post, and, when it does, the node moves into the spreader state. An illustration of the possible transition of states for a single node is provided in Figure 1. We assume (A1) that the content being shared does not express an opinion on its own, but rather the opinion is conveyed only by the person sharing it along with a comment. This assumption allows us to avoid considering the post as an external factor that can modify opinions and instead to focus on studying how the simple cascade dynamic affects opinion formation in comparison to the FJ scenario.
Figure 1: States, transitions and actions of the FJC model for a single node. The blue boxes indicate the states and the orange arrows with corresponding probabilities indicate the transitions between states. The purple arrows and boxes indicate the actions made within the corresponding states.
In our model, an individual identified as the source disseminates a piece of news to the online social network. In the rest of this paper, we streamline our discussion by initially constructing the model under the assumption that each post originates from a single seed node. However, it is worth noting that the model’s flexibility allows for straightforward extension to accommodate multiple sources, akin to the original information cascade model. In fact, including multiple sources simply affects the first level of the information cascade tree, while the rest remains unchanged. We assume (A2) that once the post is shared the first time by the seed node $r$, it spreads outwards like a tree, from the original source to the first, second, third, and subsequent in-neighbours, which we will refer to as the nodes in the first, second, third, … layer. This assumption means that the opinions of parent nodes represent the first exposure a node has in its news feed, thereby solidifying their influence. Following this initial encounter, we maintain that nodes typically become disinterested in posts shared by others. Thus, we refer to the $l$-th in-neighbourhood as the $l$-th layer (or layer $l$), and denote it by $\mathcal{O}_{l}^{(r)}$. The depth of the graph is given by the eccentricity of the root node, denoted by $\epsilon^{(r)}$, which corresponds to the maximum distance from $r$ to any other node. Accordingly, the entire tree comprising all nodes in the aware state of the cascade is represented as the sequence $\mathcal{O}^{(r)}=\left(\mathcal{O}l^{(r)}\right)_{l=0}^{\epsilon^{(r)}}$ of all layers.
A post $p$ generated at node $r$ spreads into the network over the tree $\mathcal{O}^{(r)}$. However, not all nodes in this tree receive the content, because there is a probability $\theta$ that each node will not reshare the post. The subtree of $\mathcal{O}^{(r)}$ that includes only nodes that have received the post is denoted as $\mathcal{T}^{(r,\theta)}=(\mathcal{T}^{(r,\theta)}_{l})_{l=0,\dots,\epsilon^{(r)}}$, and it holds that $\mathcal{T}^{(r,\theta)}_{l}\subseteq\mathcal{O}^{(r)}_{l}$ for each $l$. $\mathcal{T}^{(r,\theta)}$ is a stochastic realization of the tree $\mathcal{O}^{(r)}$, as indicated by the superscript $\theta$. In Figure 2, there is a simple example of a network and a corresponding propagation tree.
(a) Network
(b) $\mathcal{T}^{(1|\theta)}$
Figure 2: A toy example of a network in (a) and the corresponding propagation tree for $r=1$ in (b). Nodes 22 and $3$ update their opinion deterministically, while the others, do it with probability $\theta$.
We can describe it in formal terms in the following way. We refer to the cascade generated by the post $p$ originating in node $r$ as the collection of nodes in the tree $\mathcal{T}^{(r,\theta)}$ and we will denote it with $\mathcal{C}^{(r)}$. Please observe that nodes in layer 1 always belong to the cascade because they receive the post directly from the source, while those of layer $l\geq 2$ belong to the cascade if and only if they receive the post from nodes in layer l1l-1.
Let us indicate with $S_{i}$ the random variable modelling the event of resharing a post for a node $i\neq r$ (it takes the value 1 when the post is reshared, 0 otherwise). Then, the tree $\mathcal{T}^{(r,\theta)}$ is described as:
[…]
Please, remember that $\mathcal{N}^{-}(i)$ refers to the in-neighbours of node $i$.
For our convenience, for each node $i\in\mathcal{O}^{(r)}_{l}$ we denote with $\varphi(i)$ the set of nodes who actually influence $i$’s opinion in the stochastic realization of the cascade, i.e the nodes in layer l1l-1 that are in $i$’s out-neighbour and reshare the post. In other terms:
$\varphi(i)=\left\{j\in\mathcal{N}^{-}(i):j\in\mathcal{T}^{(r,\theta)}_{l-1}\wedge S_{j}=1\right\}$ (16)
We will refer to this set as the predecessors of $i$.
IV-B The cascade-driven Friedkin-Johnsen
We will now discuss how to integrate the independent cascade model described in Section IV-A with the FJ model from Section III-C, resulting in a FJ variant (FJC) where opinions are influenced by the sharing of posts. The FJC model operates iteratively, following specific states and transitions for each node (see Figure 1). At the start of each cascade, every node, except for the designated seed node $r$, begins in the unaware state. During each iteration, nodes $i$ that are exposed to a post from another node $s$ undergo two consecutive actions. First, these exposed nodes transition to the ”aware” state and update their opinions regarding the post. Following this, with a probability $\theta_{s,i}$, these nodes enter the spreader state, sharing the post again if they belong to a layer $l$ that is smaller than the maximum layer $\epsilon^{(r)}$. If a node does not meet this condition, the propagation of the post stops, and a new cascade is initiated. We will examine the spreading process of $N$ posts, denoted with $p_{1},\dots,p_{N}$. We make an assumption (A3) that posts propagate one at a time, meaning that simultaneous spread of two posts is not possible. While this assumption is stringent, it aids in designing the model .In future work, we will explore ways to relax this assumption and understand its impact on the model. In Section V-B, we discuss the model effectiveness in real-world settings.
Now, we want to discuss the details of the opinion updating operation. Please observe that, since we want to update opinion layer by layer, we need an asynchronous version of the FJ, similar to that of Equation (3). We focus on the spreading process of the $n$-th post $p_{n}$ generated at node $r_{n}$ which involves nodes in the cascade $\mathcal{C}^{(r_{n})}$. As the cascade starts, the opinions of nodes are equal to $\bm{u}=\bm{x}(0)$ if $n=1$, i.e. if we are considering the spreading of the first post. Otherwise, they are equal to the final opinion of nodes as the previous post $p_{n-1}$ completed spreading. Every time a node $j$ in the cascade reshares the post $p_{n}$ at a time $t$, it conveys its current opinion $x_{j}(t)$ together with the reposting. A node $i$ in layer $l$ can receive the post (which, as explained above, is a proxy for a node’s opinion) from one or more predecessors $j$. This triggers the update of node $i$’s opinion, which is carried out according to the following update rule:
$x_{i}$ $(t)=$
$=\frac{1}{|\varphi(i)|}\Bigg{\{}\sum_{j\in\varphi(i)}h_{i}\left[(1-\gamma_{ij})x_{i}(t-1)+\right$.
$\qquad\left.+\gamma_{ij}x_{j}(t-1)\right]+(1-h_{i})u_{i}\Bigg{\}},$ (17)
where $h_{i}$ and $\gamma_{i}$ are defined as in Equations 4 and 5, and $\varphi(i)$ is defined in the Equation (16). The above update rule factors in the average total influence on node $i$ by all its in-neighbours that have reshared the post in the previous layer. This means that the average is calculated only from those neighbors who have already engaged with the post and shared their opinions, thereby influencing the opinion of node $i$. Note that, differently from the asynchrounous FJ update in Equation (3), node $i$ assesses the opinions of its predecessors all at once, not in a pairwise fashion.
Differently from the FJ model in Equation (1), we update opinions node by node, following the order provided by the layer of tree $\mathcal{T}^{(r_{n},\theta)}$. Once the opinions of a single layer are updated, the post spreads again. In Algorithm 1 we summarize the entire FJC process.
Algorithm 1 FJC
[…]
4:for $r$ in $\bm{R}$ do
5:  for $l$ in $\{1,\dots,\epsilon^{(r)}\}$ do $\triangleright$ iterate over layers
6:   for $i$ in $\mathcal{T}^{(r,\theta)}_{l}$ do
7:     $x_{i}\leftarrow$ Equation  (17) $\triangleright$ update opinions
8:   end for
9:  end for
10:end for
V Performance Evaluation
In this section, we analyze the performance and validate the FJC model across different social networks. To this end, we conducted two distinct sets of experiments. The first set of experiments (Sec. V-B) focused on validating the FJC model using real-world cascade data. Here, we evaluated how well the model can approximate opinion evolution based on traces from an online social network, where assumptions (A2) and (A3) do not always hold. Specifically, (A2) posits that cascades spread in a tree-like fashion from the seed node, layer by layer, and (A3) assumes that only one post propagates at a time. Assumption (A1), however, could not be relaxed, as it underpins the model’s formulation of opinion dynamics — it assumes that the content itself is opinion-neutral, and opinions are conveyed solely by the user sharing the post. In the following, we provide a detailed description of the experimental settings and discuss the results obtained. The second set (Sec. V-C) aimed to compare the opinion dynamics produced by the FJC model with those of the classical FJ model. For this purpose, we used three social networks—two real-world and one synthetic—and generated synthetic cascades on top of them. Specifically, we selected 15 seed nodes per graph to initiate the simulated cascades.
V-A Common experimental settings
All networks used are unweighted, so the social matrix $\hat{W}$ corresponds to the adjacency matrix. As described in Section III-A, the influence matrix is derived by row-normalizing $\hat{W}$: $w_{ij}=\frac{\hat{w}{ij}}{\sum_{k}\hat{w}{ik}}$. Following our previous work [7], we considered three strategies for setting node susceptibility. The first two rely on PageRank centrality, a measure that effectively captures node influence [25]. For each node $i$ with PageRank score $\mathrm{P}_{i}$, we assigned $\lambda_{i}$ either the rescaled value of $\mathrm{P}_{i}$ or $\mathrm{P}_{i}^{-1}$, both mapped to the interval $(0,1)$, thus giving higher susceptibility to more (or less) central nodes. In the third strategy, all nodes were assigned a fixed susceptibility value of 0.6. These variations allow us to explore how susceptibility interacts with node centrality. Finally, since our primary goal is to study the impact of cascades on polarization, we initialized opinions using the polarizing vectors defined in Section III-D, ensuring a consistent and controlled setup for measuring polarization dynamics. Final opinions under cascades were then compared with those of the FJ model, whose outcomes are analytically derived via Equation (2) using the same parameters, except for the reposting probability $\theta$, which is not applicable to FJ.
V-B FJC with real cascades
V-B1 Dataset description and experimental setting
We validate the model using a real-world dataset from Bhowmick et al.  [46], which contains Twitter (now X) data related to the 2015 Nepal earthquake. This dataset is well-suited for our analysis as it includes both the follower network and tweet/retweet histories. Unlike the synthetic networks in Sec. V-C, this is a directed graph, with multiple connected components. Since FJ-based models operate within a connected structure, we extract the largest strongly connected component (LSCC). Some cascades span disconnected components, suggesting hidden links, so we restrict our analysis to cascades occurring entirely within the LSCC. The resulting graph contains 220,974 nodes, 16,772,693 edges, and 32,313 cascades. Initial opinions are set using the polarizing vectors described in Section III-D. However, their computation requires to compute of the eigensystem of an $n\times n$ dense matrix, which becomes impractical due to the extensive size of the network, even when employing optimized algorithms. To tackle this challenge, we derived a subgraph known as Nepal top cascades, which includes nodes from the 500 longest cascading events. The rationale behind focusing on the longest cascades is that they inherently present a greater challenge for our model, as they involve more complex and prolonged interactions among nodes and increase the likelihood that cascades can overlap (thus, violating assumption (A3)). Specifically, we ensured that all nodes participating in these cascades are included in the subgraph using a heuristic for the directed Steiner Tree problem [47]. Table I summarizes the key properties of this subgraph, which encompasses over 6,000 cascades.
TABLE I: Summary of the real-world Nepal dataset subgraphs used for validation
#nodes #edges #cascades max cascade length
Nepal top cascades 26321 4222064 6083 21
The dataset $D$ provides a sequence of tuples $\mathcal{C}_{D}=\big{(}(t_{1},p_{1},i_{1}),(t_{2},p_{2},i_{2}),\dots\big{)}$, where node $i_{k}$ shares post $p_{k}$ at time $t_{k}$, representing the temporal evolution of cascades. We assume that this sequence of reshares drives the opinion dynamics in the real world: each reshare is treated as an expression of opinion. At time $t_{k}$, when node $i_{k}$ shares a post, its followers update their opinions according to Equation (17). This process relies solely on assumption (A1); assumptions (A2) and (A3) are not used. We refer to this approach as real opinion dynamics. It is a data-driven heuristic and cannot be analyzed analytically.
Our goal is to evaluate how well this real dynamic is captured by the proposed FJC model (we also include the classical FJ model for reference). For the FJC model, we need to estimate the resharing probability $\theta_{i}$ for each node. Inspired by weighted cascade models [36, 48], we define $\theta_{i}$ as the ratio between the number of posts reshared by node $i$ and the number of posts it has seen from its followees. Table II shows the distribution of $\theta_{i}$, which is generally low, indicating that most nodes rarely reshare posts. Please note that these values are consistent with the retweet probabilities in [49], remaining below 0.02. Unlike the real dynamics, the FJC model processes cascades in a tree manner and sequentially. Specifically, posts are propagated from the sources to the network layer-by-layer, in accordance with assumption (A1) and each cascade starts only after the previous one ends, in accordance with assumption (A3).
TABLE II: Statistics of followers, followed and $\theta$ in the Nepal subgraph
Nepal top cascades
Min. 1st Qu. Median 3rd Qu. Max. Mean N.A.
Followers 1.0 23.0 59.0 148.0 11022.0 160.4 -
Followees 1.0 53.0 109.0 204.0 6851.0 160.4 -
$\theta$ 0.0000 0.0000 0.0000 0.0000 1.0000 0.0156 681
V-B2 Opinion evolution of real opinion dynamics vs FJC and FJ
Due to the large size of the networks, we cannot visualize the polarizing vectors, since the plots would be unreadable. Instead, our main objective with these datasets is to evaluate the robustness of the FJC model against the real opinion dynamics. We do so by comparing the distributions of opinion shifts—i.e., the differences between final and initial opinions—for the real evolution (labeled Real) and the predictions of FJC and FJ. As shown in Figure 3, FJC closely mirrors the real dynamics observed in the Nepal top cascades dataset, outperforming the standard FJ model. Please note that long cascades are more likely to violate assumption (A3), reducing FJC’s accuracy. Nevertheless, FJC remains surprisingly reliable even in this more challenging scenario.
Figure 3: ECDF of the opinion shift of nodes, i.e. final opinion minus initial opinion, for the Nepal top cascades in the three different opinion dynamics. (A) $\lambda_{i}\sim\mathrm{P}_{i}$, (B) $\lambda_{i}\sim\mathrm{P}_{i}^{-1}$, (C) $\lambda_{i}=0.6$.
V-B3 Polarization
Figure 4 illustrates the polarization outcomes across different settings. As expected, the Real opinion dynamics remain almost entirely null, since opinion shifts are negligible in all scenarios (Figure 3). Similarly, the cascade-based dynamics (FJC) exhibit very low—or even slightly negative—polarization. Both stand in sharp contrast to the classical FJ model, which consistently produces positive polarization with substantially higher magnitude. This behavior can be explained by Table II, which shows that most nodes have $\theta$ values close to zero, indicating very limited reposting activity. Consequently, the opinion dynamics remain largely inactive, preventing polarization from emerging. In Section V-C, we will further validate this hypothesis on synthetic networks, also exploring scenarios with larger $\theta$ values to assess their impact on polarization.
Figure 4: Polarization value of the different vectors in the corresponding metrics in the three different opinion dynamics. (A) $\lambda_{i}\sim\mathrm{P}_{i}$, (B) $\lambda_{i}\sim\mathrm{P}_{i}^{-1}$, (C) $\lambda_{i}=0.6$.
V-C FJC vs FJ – with synthetic cascades
V-C1 Dataset description and experimental setting
In this section, we evaluate the performance of the FJC model through simulations on three social networks: the Karate Club graph [50], a Facebook graph [51], and a synthetic Barabási-Albert network [52]. The Karate Club graph is a well-known small network with 34 nodes, while the Facebook graph includes 1519 nodes from the largest connected component of a 4039-node snapshot. The Barabási-Albert network, which captures the scale-free structure of real social networks, was generated with 100 nodes and parameter $k=6$, meaning each new node connects to 6 existing nodes. Figure 5 shows visualizations of all three graphs.
Figure 5: Networks used for the performance evaluation. In (a) and (b) are two real social networks; in (c) a synthetic network.
Since cascades are simulated, we randomly selected 15 seed nodes per network to initiate the posts, which are kept fixed across all simulations. We set the resharing probability $\theta_{s,i}$ uniformly to a single value $\theta$ for all nodes, and tested three different scenarios. In the first diffusion scenario, we set $\theta=0.01$, which is approximately the average value found in the Nepal dataset (see Table II). To explore other potential values, we referenced typical ranges established in the literature. Hodas et al. in [53] estimated that the retweet probability for each tweet seen is around 0.15% to 0.4%. Based on this, we modelled a moderate diffusion scenario with $\theta=0.1$, and for maximum diffusion, we used $\theta=0.5$. For each combination of network, susceptibility setting, reposting probability $\theta$, and initial opinion values, we performed 1000 independent simulation runs.
V-C2 Opinion evolution under FJC vs FJ
We begin by analyzing how cascades influence final opinions under the FJC model compared to the classical FJ model. Due to space constraints, we will present just $\bm{u}_{B_{1}(1)}$ and $\bm{u}_{V_{>1}}^{heu}$, the best local sub-optimal polarizing vectors that maximise $P_{4}$ and $P_{2}-P_{3}$ polarization, respectively. Please refer to the Supplemental Information (SI) for the full results. In Figures 6-8, each arrow represents a node’s opinion shift—from its initial opinion (prejudice) to its final FJC opinion. The black dot indicates the corresponding final opinion under the FJ model. Panels A and B show cases where susceptibility is proportional or inversely proportional to node centrality, respectively; panel C shows uniform susceptibility ($\lambda=0.6$).
We begin our discussion with Figure 8, which corresponds to the scenario with the highest propagation ($\theta=0.5$), as this setting highlights the effects of cascades on opinion dynamics at a larger scale. Afterwards, we will examine how these trends change when $\theta$ is smaller. Let us first consider panel C, where all nodes share the same susceptibility. Compared to the classical FJ model, highly central nodes in the FJC model tend to retain their initial opinions more strongly—evident from the shorter arrows relative to the distance to the FJ outcomes (black dots). This leads to a wider spread of final opinions among central nodes. In contrast, peripheral nodes are more easily influenced, and their opinions shift closer to those of the central ones. Overall, cascades amplify the role of network topology: central nodes act as opinion anchors, while peripheral nodes become more susceptible followers. When susceptibility is inversely proportional to centrality (panel B), central nodes are already stubborn by design in both models. Final opinions are mostly similar across all networks, although we can observe slightly smaller opinion shifts in FJC. This reduction is due to the parameter $\theta$, which limits opinion propagation. This trend is further confirmed by the figures in the SI, where we also tested even larger values for this parameter: as $\theta$ increases, peripheral nodes shift more strongly in the FJC model, becoming increasingly influenced by central nodes. Finally, with susceptibility proportional to centrality (panel A), network topology and susceptibility act in opposing directions. Peripheral nodes are nearly immobile due to low susceptibility, while central nodes—despite being more susceptible—change little due to their structural position. As a result, final opinions under FJC remain close to initial ones, more so than in the FJ model. This highlights how, in cascade-driven dynamics, centrality outweighs susceptibility, with central nodes maintaining influence regardless of the surrounding opinion mass.
Similar behaviors, though less pronounced, are observed with smaller values of $\theta$ ($\theta=0.01,0.1$) in Figures 6 and 7. In general, peripheral nodes exhibit reduced opinion shifts as $\theta$ decreases, reflecting their stronger dependence on the probability of content being reshared. Notably, in the Karate network, results remain almost unchanged across different $\theta$ values, as its small size allows most nodes to be reached by the cascade even when propagation is limited. Interestingly, central nodes in the FJC model maintain stable behavior even with tighter propagation. This resilience is an intrinsic property of the model, stemming from the structured, sequential nature of its cascades. In fact, highly connected nodes are typically exposed to content early, when active sources are few, making them less susceptible to external influence.
$\bm{\theta=0.01}$
Figure 6: Polarizing opinions vs. node centrality with $\theta=0.01$. Each arrow represents a node, starting at its initial opinion (prejudice) and ending at its final opinion under the FJC model. Arrow color indicates node susceptibility; black dots show the final opinions under the FJ model. (A) $\lambda_{i}\propto\mathrm{P}_{i}$; (B) $\lambda_{i}\propto\mathrm{P}_{i}^{-1}$; (C) $\lambda_{i}=0.6$. Dot-dashed and dashed lines show the average initial and final FJC opinions.
$\bm{\theta=0.1}$
Figure 7: Polarizing opinions vs. node centrality with $\theta=0.1$. Each arrow represents a node, starting at its initial opinion (prejudice) and ending at its final opinion under the FJC model. Arrow color indicates node susceptibility; black dots show the final opinions under the FJ model. (A) $\lambda_{i}\propto\mathrm{P}_{i}$; (B) $\lambda_{i}\propto\mathrm{P}_{i}^{-1}$; (C) $\lambda_{i}=0.6$. Dot-dashed and dashed lines show the average initial and final FJC opinions.
$\bm{\theta=0.5}$
Figure 8: Polarizing opinions vs. node centrality with $\theta=0.5$. Each arrow represents a node, starting at its initial opinion (prejudice) and ending at its final opinion under the FJC model. Arrow color indicates node susceptibility; black dots show the final opinions under the FJ model. (A) $\lambda_{i}\propto\mathrm{P}_{i}$; (B) $\lambda_{i}\propto\mathrm{P}_{i}^{-1}$; (C) $\lambda_{i}=0.6$. Dot-dashed and dashed lines show the average initial and final FJC opinions.
V-C3 Polarization
Figure 9 shows how opinion dynamics under FJC affect polarization compared to the FJ model, measured as the change in polarization (Equation (14)) between final and initial opinions. We report results for each susceptibility setting—proportional (A), inversely proportional (B), and uniform (C)—and for each $\theta$ value (differently colored bars for FJC). We show the polarization metric maximized by each polarizing vector in the FJ model: $P_{4}$ for $\bm{u}_{B_{1}(1)}$ and $P_{2}$ for the others. We first consider FJC with $\theta=0.5$ (purple bars). When susceptibility is uniform (panel C), FJC consistently yields higher polarization than FJ, as central nodes remain influential and pull others toward their opinions. A similar, though less pronounced, pattern appears with inverse-centrality susceptibility (panel B), since less central nodes become more vulnerable under FJC. An exception arises for $\bm{u}_{{V_{>1}}^{heu}}$ in the BA network, where polarization decreases because the parameter $\theta$ limits opinion propagation, as discussed in the previous section. Interestingly, when susceptibility is proportional to centrality (panel A), FJC often reduces polarization compared to FJ. Here, central nodes are nearly immovable, and peripheral nodes, unable to influence them, shift closer to reduce cognitive dissonance—flattening the opinion landscape. An exception occurs in the Facebook network with $\bm{u}_{B_{1}(1)}$, where the initial opinion aligns with a central node that remains firm, attracting others and increasing polarization. For $\theta=0.01$ and $\theta=0.1$ (green and blue bars), the cascade effect is weak due to limited reposting, resulting in polarization trends similar to those at higher $\theta$ but with lower magnitude. In some cases, the effect even reverses, leading to depolarization—primarily driven by the few nodes directly connected to the sources, which are the only ones likely to update their opinions. For instance, in the Karate network with proportional susceptibility and the vector $\bm{u}_{B_{1}(1)}$, most neighbors of the sources have low susceptibility and thus remain unchanged. The only significant shift comes from the source node itself, which lowers its opinion from 1, reducing overall polarization as measured by $P_{4}$.
Figure 9: Polarization of FJ and FJC (with different $\theta$ values) for (A) $\lambda_{i}\sim\mathrm{P}_{i}$, (B) $\lambda_{i}\sim\mathrm{P}_{i}^{-1}$, (C) $\lambda_{i}=0.6$.
VI Conclusions
This study explores the dynamics of opinion formation and dissemination within online social networks, particularly under the influence of information cascades. Specifically, the paper proposes the Friedkin Johnsen on Cascade model (FJC), which integrates opinion dynamics and information cascades. To the best of our knowledge, this is the first model that allows for the investigation of how content dissemination can affect opinion formation. We validated the FJC model, utilising a real cascade dataset. Our results demonstrate that the FJC model can reliably approximate final opinions accurately, even when many of the underlying assumptions of the model are relaxed or only partially met. We evaluate the model using both synthetic and real social networks and highlight that the cascades reinforce the strength of central nodes in influencing opinions. These central nodes tend to be less vulnerable to opposing opinions, making it unlikely for them to change their views even when pressured by a large majority. This occurs when content is free to circulate within the network. Conversely, when participation is low, content cannot go viral, disrupting the opinion dynamics. The model builds upon certain assumptions regarding content dissemination and opinion evolution.